OpenStudy (anonymous):

sam deposited Rs. 350 per month for 6 years in a recurring deposit account in a post office. If she received Rs.32,865, then what is the rate of interest?

OpenStudy (anonymous):

You can use Recurring Deposit Maturity Value Formula to solve this: \[M = \frac{ R [(1+r)^{n} - 1 ] }{ (1-(1+r)^{-1/3}) }\] Where, M = Maturity value R = Monthly Installment r = Rate of Interest (i) / 400 n = Number of Quarters At the end of every month, you put Rupees 350 in to recurring deposit that pay r%, compounded monthly. So, the Future Value after 6 years (72Months) is Rs.32,865 Then, M = 32,865; i=r%/12; N = 72; R = 350 By substituting these values in RD Formula u can find the rate of interest.